OpenStudy (anonymous):

If n = 8 and 16 · 2m = 4n – 8, then m = ?

OpenStudy (anonymous):

Substitute n=8 into the equation; 16*2m = 4*8-8 which gives 32m = 24 then divide both sides by 32 m=3/4

OpenStudy (anonymous):

I don't understand.

OpenStudy (imtiaz7):

16.2m=4n-8 32m=4n-8 therefore n=8 by putting the value of n in above equation: 32m=4(8)-8 32m=32-8 32m=24 m=24/32 m=6/8 or m=3/4

OpenStudy (anonymous):

Which part is confusing you?

OpenStudy (anonymous):

The part where you substitute n. how is it not an exponent anymore?

OpenStudy (anonymous):

Where does it say that it's an exponent? Does the question say 4 to the power of n?

OpenStudy (anonymous):

yeah. oh crap. the question is messed up here. it didn't paste properly.

OpenStudy (anonymous):

if n = 8 and 16 * 2^m = 4^ (n-8), them m = ?

OpenStudy (anonymous):

i'm sorry about that.

OpenStudy (anonymous):

Ok no problem. Anything to the power of zero equals 1. Does this help?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

So now you have 2^m = 1/16, can you finish it?

OpenStudy (anonymous):

give me a sec please =)

OpenStudy (anonymous):

nope. how does 1 apply here?

OpenStudy (anonymous):

Okay there's a couple of ways to do this now, if you've been taught logarithms in school you could take log_2 (logarithm base 2) of both sides to get m = log_2(1/16)

OpenStudy (anonymous):

i haven't worked on logs yet. should i get to it first before i solve these problems? =)

OpenStudy (anonymous):

Or if you just think about it 2^(-1) =1/2, and 1/2 squared is 1/4, then 1/4 squared is 1/16, so by inspection you can see that 2^(-4) = 1/16

OpenStudy (anonymous):

But yeah for more complex problems you would need logs

OpenStudy (anonymous):

i'll get to it then. thank you!